# On pro-cdh descent on derived schemes

### v1

@@ 3. Modifications of derived schemes
Assume that $X$ carries an ample line bundle. Let $U\subseteq X$ be a quasi-compact open subscheme. Let $f\colon Y\to X$ be an lafp $U$-modification of $X$. Then there exists a commutative diagram
@@
where $g$ is an lafp $U$-modification of $Y$, $h$ is an lafp closed $U$-modification, and $p$ is a derived blowup with center set-theoretically contained in $X\setminus U$. Moreover, $Y^{\prime}$ carries a $(p\circ h)$-ample line bundle.
@@
As $f\colon Y\to X$ is proper and lafp, the morphism of classical schemes $\prescript{\mathrm{cl}}{}{f}\colon\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is proper and classically locally of finite presentation, hence classically of finite presentation. It is also an isomorphism over $\prescript{\mathrm{cl}}{}{U}$. By [RG71, Cor. 5.7.12] (or [Sta23, Tag 081T]), there exists a $\prescript{\mathrm{cl}}{}{U}$-admissible blowup333i.e., a blowup in a finitely presented closed subscheme of $\prescript{\mathrm{cl}}{}{X}$ which is set-theoretically contained in the complement of $\prescript{\mathrm{cl}}{}{U}\subseteq\prescript{\mathrm{cl}}{}{X}$ $Y_{0}\to\prescript{\mathrm{cl}}{}{X}$ such that the morphism $Y_{0}\to\prescript{\mathrm{cl}}{}{X}$ factors through $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$. For the constructions to come, we need the open immersion $\prescript{\mathrm{cl}}{}{U}\to Y_{0}$ to be affine. As this is not necessarily the case, we make a further blowup: As $\prescript{\mathrm{cl}}{}{U}$ is quasi-compact, there exists a finitely presented closed subscheme $T_{0}\hookrightarrow Y_{0}$ whose underlying topological space is $Y_{0}\setminus\prescript{\mathrm{cl}}{}{U}$ [^4] For example, one can use absolute Noetherian approximation [TT90, Thm. C.9, C.2] to see this. Let $Y_{1}\to Y_{0}$ be the blowup of $Y_{0}$ in $T_{0}$. Then the canonical open immersion $\prescript{\mathrm{cl}}{}{U}\to Y_{1}$ is affine as its complement is a Cartier divisor. As the composite of two $\prescript{\mathrm{cl}}{}{U}$-admissible blowups is a $\prescript{\mathrm{cl}}{}{U}$-admissible blowup [Sta23, Tag 080L], the composition $Y_{1}\to Y_{0}\to\prescript{\mathrm{cl}}{}{X}$ is a $\prescript{\mathrm{cl}}{}{U}$-admissible blowup, say $Y_{1}=\operatorname{Bl}_{S_{0}}(\prescript{\mathrm{cl}}{}{X})$ for some classically finitely presented closed subscheme $S_{0}\hookrightarrow\prescript{\mathrm{cl}}{}{X}$.
@@
As $X$ carries an ample line bundle, Lemma 2.6 implies the existence of a quasi-smooth closed immersion of derived schemes $S\hookrightarrow X$ whose classical truncation is $S_{0}\hookrightarrow\prescript{\mathrm{cl}}{}{X}$. Let $p\colon\widetilde{X}\to X$ be the derived blowup of $X$ in $S$. Then there is a canonical closed immersion $Y_{1}\hookrightarrow\prescript{\mathrm{cl}}{}{\widetilde{X}}$, which is an isomorphism over $\prescript{\mathrm{cl}}{}{U}$. Note that $Y_{1}$ need not be classically of finite presentation over $X$. However, by [GD71, Cor. 6.9.15] we can write the ideal sheaf defining $Y_{1}$ in $\prescript{\mathrm{cl}}{}{\widetilde{X}}$ as a filtered colimit of sub-ideal sheaves of finite type. Passing to relative spectra, we deduce that the closed immersion $Y_{1}\hookrightarrow\prescript{\mathrm{cl}}{}{\widetilde{X}}$ can be written as a cofiltered limit of classically finitely presented closed immersions $Y_{\alpha}\to\prescript{\mathrm{cl}}{}{\widetilde{X}}$ all of which are isomorphisms over $\prescript{\mathrm{cl}}{}{U}$. As $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is classically of finite presentation, there exists an $\alpha$ such that the $\prescript{\mathrm{cl}}{}{X}$-morphism $Y_{1}\to\prescript{\mathrm{cl}}{}{Y}$ factors through a morphism $Y_{\alpha}\to\prescript{\mathrm{cl}}{}{Y}$, see [Gro66, Prop. 8.14.2]. Thus, so far, we have constructed a commutative diagram of classical schemes finitely presented over $\prescript{\mathrm{cl}}{}{X}$
@@
Note that $h_{2}$ is a closed immersion, as this only depends on the underlying map of classical schemes. Moreover, all morphisms in the above diagram are $U$-modifications. However, $h_{2}$ and $g_{2}$ need not be locally of almost finite presentation. In order to remedy this, we consider the induced morphism $k_{2}=(h_{2},g_{2})\colon Y_{2}\to\widetilde{X}\times_{X}Y$. As $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is separated, and as $h_{2}$ is a closed immersion, $k_{2}$ is a closed immersion, too. It is also an isomorphism over $U$. By construction, the map of underlying classical schemes $\prescript{\mathrm{cl}}{}{k_{2}}\colon Y_{\alpha}=\prescript{\mathrm{cl}}{}{Y}_{2}\to\prescript{\mathrm{cl}}{}{(}\widetilde{X}\times_{X}Y)$ is classically of finite presentation. We can hence apply Lemma 2.2 to obtain a factorization of $k_{2}$ through a closed derived subscheme $k^{\prime}\colon Y^{\prime}\hookrightarrow\widetilde{X}\times_{X}Y$ such that $k^{\prime}$ is locally of finite presentation and an isomorphism over $U$, and $\prescript{\mathrm{cl}}{}{Y}^{\prime}\cong\prescript{\mathrm{cl}}{}{Y}_{2}$. The composite $h\colon Y^{\prime}\hookrightarrow\widetilde{X}\times_{X}Y\to\widetilde{X}$ is then lafp and an isomorphism over $U$. As its underlying map of classical schemes identifies with $\prescript{\mathrm{cl}}{}{h}_{2}\colon\prescript{\mathrm{cl}}{}{Y}_{2}\hookrightarrow\prescript{\mathrm{cl}}{}{\widetilde{X}}$, it is a closed immersion. Similarly, the composite $g\colon Y^{\prime}\to\widetilde{X}\times_{X}Y\to Y$ is lafp and an isomorphism over $U$. This finishes the construction of the asserted commutative diagram.
@@ 4. Pro-cdh descent for connective localizing invariants
In this section, we prove our main results on pro-descent for localizing invariants. The strategy is the same as in [KST18b]: We first prove the result for the special cases of derived blowups and finite modifications and then use the geometric input from Theorem 3.2 to handle the general case.
@@
A version of the following Proposition was first proven in [KST18b]. There all schemes were assumed to be classical Noetherian schemes, $E$ was $K$-theory, and the result only gave a weakly cartesian square of pro-spectra, see below for this notion. In a letter to Kerz, Antieau [Ant18] described a simplification of the proof which at the same time gives a cartesian square. We thank Ben Antieau for allowing us to include his argument in our paper.
@@
We show below that this system is weakly contractible. We then get weak equivalences $J\xrightarrow{\simeq}\{J/\!\!/f_{1}^{\alpha}\}_{\alpha}$ and $J\xrightarrow{\simeq}\{J/\!\!/f_{2}^{\alpha}\}_{\alpha}\xrightarrow{\simeq}\{J/\!\!/f_{1}^{\alpha},f_{2}^{\alpha}\}_{\alpha}$, and so on, so {(7)}$\}_{\alpha}$ is indeed weakly cartesian.
@@
As a special case, if the map $f$ in Proposition 4.2 is a closed immersion of classical qcqs schemes which is classically finitely presented, then the conclusion of the proposition holds, i.e. (6) is weakly cartesian.
@@
If $X$ and $Y$ are Noetherian classical schemes, and $f$ is classically of finite type, then $f$ is lafp (see 2.4). Moreover, the formal derived completions are equivalent to the classical formal completions (see 2.5). We thus recover the classical pro-cdh descent statement as formulated (for $K$-theory) for example in [KST18b, Thm. A].
@@ 5. Pro-cdh descent for the cotangent complex and motivic cohomology
The following theorem was suggested by Matthew Morrow. It generalizes [Mor16a, Thm. 2.4] (see also [EM23, Lemma 8.5]).
@@
6. Generalised Weibel vanishing
In this section we prove our generalised Weibel vanishing result. As its formulation involves the valuative dimension, we start by recalling the latter. The valuative dimension of a commutative ring was introduced and studied by Jaffard [Jaf60]; we refer to [EHIK20, §2.3] for an account. For an integral domain $A$, it is defined as
@@ 6. Generalised Weibel vanishing
We always have $\dim(X)\leq\operatorname{vdim}(X)$ and equality holds if $X$ is Noetherian [Jaf60, Ch. IV, Thm. 1, Cor. 2 of Thm. 5]. Note that the valuative dimension only depends on the underlying reduced scheme of a scheme. In particular, it makes sense to talk about the valuative dimension of a closed subset of a scheme, by equipping it with any closed subscheme structure.
@@
Let now $X$ be a derived scheme. The Krull dimension $\dim(X)$ of $X$ is defined to be the Krull dimension of the underlying topological space of $X$. The valuative dimension $\operatorname{vdim}(X)$ is defined to be the valuative dimension of the underlying classical scheme $\prescript{\mathrm{cl}}{}{X}$.

### v2

@@ 3. Modifications of derived schemes
Assume that $X$ carries an ample line bundle. Let $U\subseteq X$ be a quasi-compact open subscheme. Let $f\colon Y\to X$ be a $U$-modification of $X$, whose underlying map of classical schemes $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is classically finitely presented. Then there exists a commutative diagram
@@
where $g$ is $U$-modification of $Y$, $h$ is a closed $U$-modification, $p$ is a derived blowup with center set-theoretically contained in $X\setminus U$, and the induced map $(h,g)\colon Y^{\prime}\to\widetilde{X}\times_{X}Y$ is locally of finite presentation. In particular, $g$ is locally of finite presentation, and, if $f$ is locally of finite presentation (or lafp, or locally of finite generation to order $n$, respectively), then $h$ is locally of finite presentation (or lafp, or locally of finite generation to order $n$).
Moreover, $Y^{\prime}$ carries a $(p\circ h)$-ample line bundle.
@@
By assumption, the morphism of classical schemes $\prescript{\mathrm{cl}}{}{f}\colon\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is proper and classically of finite presentation. It is also an isomorphism over $\prescript{\mathrm{cl}}{}{U}$. By [RG71, Cor. 5.7.12] (or [Sta25, Tag 081T]), there exists a $\prescript{\mathrm{cl}}{}{U}$-admissible blowup333i.e., a blowup in a finitely presented closed subscheme of $\prescript{\mathrm{cl}}{}{X}$ which is set-theoretically contained in the complement of $\prescript{\mathrm{cl}}{}{U}\subseteq\prescript{\mathrm{cl}}{}{X}$ $Y_{0}\to\prescript{\mathrm{cl}}{}{X}$ such that the morphism $Y_{0}\to\prescript{\mathrm{cl}}{}{X}$ factors through $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$. For the constructions to come, we need the open immersion $\prescript{\mathrm{cl}}{}{U}\to Y_{0}$ to be affine. As this is not necessarily the case, we make a further blowup: As $\prescript{\mathrm{cl}}{}{U}$ is quasi-compact, there exists a finitely presented closed subscheme $T_{0}\hookrightarrow Y_{0}$ whose underlying topological space is $Y_{0}\setminus\prescript{\mathrm{cl}}{}{U}$: This follows for instance from [GD71, Cor. 6.9.15] which allows us to write the ideal sheaf of some closed subscheme of $Y_{0}$ with support $Y_{0}\setminus\prescript{\mathrm{cl}}{}{U}$ as a filtered colimit of quasi-coherent sub-ideal sheaves of finite type. By quasi-compactness of $\prescript{\mathrm{cl}}{}{U}$, the closed subscheme defined by one of them will have support $Y_{0}\setminus\prescript{\mathrm{cl}}{}{U}$. Let $Y_{1}\to Y_{0}$ be the blowup of $Y_{0}$ in $T_{0}$. Then the canonical open immersion $\prescript{\mathrm{cl}}{}{U}\to Y_{1}$ is affine as its complement is a Cartier divisor. As the composite of two $\prescript{\mathrm{cl}}{}{U}$-admissible blowups is a $\prescript{\mathrm{cl}}{}{U}$-admissible blowup [Sta25, Tag 080L], the composition $Y_{1}\to Y_{0}\to\prescript{\mathrm{cl}}{}{X}$ is a $\prescript{\mathrm{cl}}{}{U}$-admissible blowup, say $Y_{1}=\operatorname{Bl}_{S_{0}}(\prescript{\mathrm{cl}}{}{X})$ for some classically finitely presented closed subscheme $S_{0}\hookrightarrow\prescript{\mathrm{cl}}{}{X}$.
@@
As $X$ carries an ample line bundle, Lemma 2.6 implies the existence of a quasi-smooth closed immersion of derived schemes $S\hookrightarrow X$ whose classical truncation is $S_{0}\hookrightarrow\prescript{\mathrm{cl}}{}{X}$. Let $p\colon\widetilde{X}\to X$ be the derived blowup of $X$ in $S$. Then there is a canonical closed immersion $Y_{1}\hookrightarrow\prescript{\mathrm{cl}}{}{\widetilde{X}}$, which is an isomorphism over $\prescript{\mathrm{cl}}{}{U}$. Note that $Y_{1}$ need not be classically of finite presentation over $X$. However, using [GD71, Cor. 6.9.15] again we see that the closed immersion $Y_{1}\hookrightarrow\prescript{\mathrm{cl}}{}{\widetilde{X}}$ can be written as a cofiltered limit of classically finitely presented closed immersions $Y_{\alpha}\to\prescript{\mathrm{cl}}{}{\widetilde{X}}$ all of which are isomorphisms over $\prescript{\mathrm{cl}}{}{U}$. As $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is classically of finite presentation, there exists an $\alpha$ such that the $\prescript{\mathrm{cl}}{}{X}$-morphism $Y_{1}\to\prescript{\mathrm{cl}}{}{Y}$ factors through a morphism $Y_{\alpha}\to\prescript{\mathrm{cl}}{}{Y}$, see [Gro66, Prop. 8.14.2]. Thus, so far, we have constructed a commutative diagram of classical schemes finitely presented over $\prescript{\mathrm{cl}}{}{X}$
@@
Note that $h_{2}$ is a closed immersion, as this only depends on the underlying map of classical schemes. Moreover, all morphisms in the above diagram are $U$-modifications. However, $h_{2}$ and $g_{2}$ need not satisfy the desired finiteness condition. In order to remedy this, we consider the induced morphism $k_{2}=(h_{2},g_{2})\colon Y_{2}\to\widetilde{X}\times_{X}Y$. As $\prescript{\mathrm{cl}}{}{Y}\to\prescript{\mathrm{cl}}{}{X}$ is separated, and as $h_{2}$ is a closed immersion, $k_{2}$ is a closed immersion, too. It is also an isomorphism over $U$. By construction, the map of underlying classical schemes $\prescript{\mathrm{cl}}{}{k_{2}}\colon Y_{\alpha}=\prescript{\mathrm{cl}}{}{Y}_{2}\to\prescript{\mathrm{cl}}{}{(}\widetilde{X}\times_{X}Y)$ is classically of finite presentation. We can hence apply Lemma 2.2 to obtain a factorization of $k_{2}$ through a closed derived subscheme $k^{\prime}\colon Y^{\prime}\hookrightarrow\widetilde{X}\times_{X}Y$ such that $k^{\prime}$ is locally of finite presentation and an isomorphism over $U$, and $\prescript{\mathrm{cl}}{}{Y}^{\prime}\cong\prescript{\mathrm{cl}}{}{Y}_{2}$. Define $g$ and $h$ to be the composites of $k^{\prime}$ with the two projections from $\widetilde{X}\times_{X}Y$ to $Y$ and $\widetilde{X}$, respectively. Both are $U$-modifications. As $p\colon\widetilde{X}\to X$ is locally of finite presentation, so is the first projection, and hence $g$. If $f$ is locally of finite presentation (or lafp, or locally of finite generation to order $n$), then so is the second projection and hence also $h$. As the underlying map of classical schemes $\prescript{\mathrm{cl}}{}{h}$ identifies with $\prescript{\mathrm{cl}}{}{h}_{2}\colon\prescript{\mathrm{cl}}{}{Y}_{2}\hookrightarrow\prescript{\mathrm{cl}}{}{\widetilde{X}}$, it is a closed immersion. This finishes the construction of the asserted commutative diagram, and the proof of the required finiteness conditions.
@@ 4. Pro-cdh descent for connective localizing invariants
In this section, we prove our main results on pro-descent for localizing invariants. The strategy is the same as in [KST18]: We first prove the result for the special cases of derived blowups and finite modifications and then use the geometric input from Theorem 3.2 to handle the general case.
@@
A version of the following Proposition was first proven in [KST18]. There all schemes were assumed to be classical Noetherian schemes, $E$ was $K$-theory, and the result only gave a weakly cartesian square of pro-spectra, see below for this notion. In a letter to Kerz, Antieau [Ant18] described a simplification of the proof which at the same time gives a cartesian square. We thank Ben Antieau for allowing us to include his argument in our paper.
@@
We show below that this system is weakly contractible. We then get weak equivalences $J\xrightarrow{\simeq}\{J/\!\!/f_{1}^{\alpha}\}_{\alpha}$ and, taking derived quotients by powers of $f_{2}$, $\{J/\!\!/f_{2}^{\alpha}\}_{\alpha}\xrightarrow{\simeq}\{J/\!\!/f_{1}^{\alpha},f_{2}^{\alpha}\}_{\alpha}$. Composing with $J\xrightarrow{\simeq}\{J/\!\!/f_{2}^{\alpha}\}_{\alpha}$ we get the weak equivalence $J\xrightarrow{\simeq}\{J/\!\!/f_{1}^{\alpha},f_{2}^{\alpha}\}_{\alpha}$ and continuing like this we finally arrive at the weak equivalence $J\xrightarrow{\simeq}\{J/\!\!/f_{1}^{\alpha},\dots,f_{r}^{\alpha}\}_{\alpha}$. So {(7)}$\}_{\alpha}$ is indeed weakly cartesian.
@@
As a special case, if the map $f$ in Proposition 4.2 is a closed immersion of classical qcqs schemes which is classically finitely presented, then the conclusion of the proposition holds, i.e. (6) is weakly cartesian. On the other hand, we cannot drop the finite presentation assumption, as the following example shows.
Example 4.4.
Let $A$ be a discrete, commutative ring containing an element $f$ and a non-trivial ideal $J$ such that $J[f^{-1}]=0$ and $J$ is $f$-divisible. Note that this implies that $J^{2}=0$. For example, take $A=\mathbb{Z}\oplus\mathbb{Q}_{p}/\mathbb{Z}_{p}$, $f=p$, $J=\mathbb{Q}_{p}/\mathbb{Z}_{p}$.
The closed immersion $Y=\operatorname{Spec}(A/J)\to X=\operatorname{Spec}(A)$ is then an isomorphism outside $Z=V(f)$. We claim that the square
(9) ${K(X)}$${K(X_{Z}^{\wedge})}$${K(Y)}$${K(Y_{Z}^{\wedge})}$
is not weakly cartesian.
Aiming at a contradiction, assume (9) is weakly cartesian. Fix some positive integer $n$. As $J$ is $f$-divisible,
$J/\!\!/f^{n}=\operatorname{fib}(A/\!\!/f^{n}\to(A/J)/\!\!/f^{n})$
is 1-connective. As a consequence, the map of $K$-theory spectra
$K(A/\!\!/f^{n})\to K((A/J)/\!\!/f^{n})$
is 2-connective. So also the right vertical map, and hence the left vertical map in (9) is 2-connective. In particular, the map $K_{1}(A)\to K_{1}(A/J)$ is an isomorphism. However, its retract $A^{*}\to(A/J)^{*}$ is not injective as every $1+x$ with $x\in J$ is an element of its kernel.
@@
If $X$ and $Y$ are Noetherian classical schemes, and $f$ is classically of finite type, then $f$ is lafp (see 2.4). Moreover, the formal derived completions are equivalent to the classical formal completions (see 2.5). We thus recover the classical pro-cdh descent statement as formulated (for $K$-theory) for example in [KST18, Thm. A].
@@ 5. Pro-cdh descent for the cotangent complex and motivic cohomology
The following theorem was suggested by Matthew Morrow. It generalizes [Mor16b, Thm. 2.4] (see also [EM23, Lemma 8.5]).
@@
6. Generalized Weibel vanishing
In this section we prove our generalized Weibel vanishing result. As its formulation involves the valuative dimension, we start by recalling the latter (Subsection 6.1). In Subsection 6.2 we generalize some results about reductions of ideals from [HS06] to non-Noetherian rings. These will go into the proof of Weibel vanishing in Subsection 6.3.
6.1. Valuative dimension
The valuative dimension of a commutative ring was introduced and studied by Jaffard [Jaf60]; we refer to [EHIK20, §2.3] for an account and to [KS24, Lemma 7.2] which reproves some of Jaffard’s key results in modern language. For an integral domain $A$, it is defined as
@@ 6.1. Valuative dimension
We remark that that $\operatorname{vdim}(X)=\sup\dim\operatorname{RZ}(X_{\lambda})$ is the supremum of the Krull dimensions of the Riemann–Zariski spaces as $X_{\lambda}$ ranges over the irreducible components of $X$. That is, the supremum of the Krull dimensions of all blowups (of finite type) of all $X_{\lambda}$. We always have $\dim(X)\leq\operatorname{vdim}(X)$ and equality holds if $X$ is Noetherian [Jaf60, Ch. IV, Thm. 1, Cor. 2 of Thm. 5]. Note that the valuative dimension only depends on the underlying reduced scheme of a scheme. In particular, it makes sense to talk about the valuative dimension of a closed subset of a scheme, by equipping it with any closed subscheme structure.
@@
Let now $X$ be a derived or spectral scheme. The Krull dimension $\dim(X)$ of $X$ is defined to be the Krull dimension of the underlying topological space of $X$. The valuative dimension $\operatorname{vdim}(X)$ is defined to be the valuative dimension of the underlying classical scheme $\prescript{\mathrm{cl}}{}{X}$.
We will need the following result.
Lemma 6.1.
For a scheme $X$ and a point $x\in X$ with closure $\overline{\{x\}}$ in $X$,
In particular,
Proof.
By definition it suffices to prove the lemma in case $X=\operatorname{Spec}(A)$ with $A$ integral. In this case, the first assertion is [Jaf60, Ch. IV, Prop. 2]; indeed, this follows from the basic facts that for any field extension $L/K$ any valuation $R\subseteq K$ admits at least one extension $S\subseteq L$, and $\operatorname{Spec}(S)\to\operatorname{Spec}(R)$ is always surjective. The second assertion follows from the fact that $\dim(Y)\leq\operatorname{vdim}(Y)$ for every scheme $Y$. ∎
6.2. Reductions
Let $R$ be a ring, and let $J\subseteq I$ be ideals in $R$. Then $J$ is called a reduction of $I$ if there is a positive integer $n$ such that $JI^{n}=I^{n+1}$. Associated with an ideal $I$ we have its Rees algebra
$R[It]=\sum_{n\geq 0}I^{n}t^{n}\subseteq R[t],\quad R[It]\cong\bigoplus_{n\geq 0}I^{n}$
and its extended Rees algebra
$R[It,t^{-1}]=\sum_{n<0}Rt^{n}+\sum_{n\geq 0}I^{n}t^{n}\subseteq R[t,t^{-1}].$
So $\operatorname{Proj}(R[It])$ is the blowup of $\operatorname{Spec}(R)$ in $\operatorname{Spec}(R/I)$, and $\operatorname{Spec}(R[It,t^{-1}])$ is a flat deformation of $\operatorname{Spec}(R)$ to the normal cone $\operatorname{Spec}(\bigoplus_{n\geq 0}I^{n}/I^{n+1})$.
Lemma 6.2.
Assume that $J\subseteq I$ are ideals in $R$.
  1. (1)
    If $I$ is finitely generated, and $J$ is a reduction of $I$, then the ring extension $R[Jt]\subseteq R[It]$ is finite.
  2. (2)
    Conversely, if the ring extension $R[Jt]\subseteq R[It]$ is finite, then $J$ is a reduction of $I$.
Proof.
This is [HS06, Th. 8.2.1] which is formulated for ideals in Noetherian rings. The proof given there works verbatim in the asserted generality. ∎
Let $R$ be a local ring with maximal ideal $\mathfrak{m}$ and residue field $\kappa$. Let $I\subset R$ be a finitely generated, proper ideal. Then the fibre cone
$\mathcal{F}_{I}(R)=R[It]\otimes^{\heartsuit}_{R}\kappa=\kappa\oplus I/\mathfrak{m}I\oplus I^{2}/\mathfrak{m}I^{2}\oplus\dots$
is a finitely generated $\kappa$-algebra and in particular Noetherian. Its Krull dimension $\ell(I):=\dim\mathcal{F}_{I}(R)$ is called the analytic spread of $I$.
Proposition 6.3.
Let $R$ be a local ring, $I\subseteq R$ a finitely generated, proper ideal. Then the analytic spread of $I$ is bounded by the valuative dimension of $R$:
$\ell(I)\leq\operatorname{vdim}(R).$
Proof.
This proof follows [HS06, Prop. 5.1.6], which treats the Noetherian case. We may assume that $\operatorname{vdim}(R)$ is finite. Let $\operatorname{gr}_{I}(R)=R[It]\otimes^{\heartsuit}_{R}R/I=R/I\oplus I/I^{2}\oplus\dots=R[It,t^{-1}]/(t^{-1})$. As $\mathcal{F}_{I}(R)$ is a quotient of $\operatorname{gr}_{I}(R)$ and $\mathcal{F}_{I}(R)$ is Noetherian, we have
$\ell(I)=\operatorname{vdim}\mathcal{F}_{I}(R)\leq\operatorname{vdim}\operatorname{gr}_{I}(R).$
As $t^{-1}$ is a non-zero divisor in $R[It,t^{-1}]$, it is not contained in any minimal prime ideal of $R[It,t^{-1}]$. Thus $\operatorname{vdim}\operatorname{gr}_{I}(R)\leq\operatorname{vdim}R[It,t^{-1}]-1$ by [EHIK20, Prop. 2.3.2(4)]. We claim that $\operatorname{vdim}R[It,t^{-1}]=\operatorname{vdim}R+1$. This will finish the proof. To prove the claim, as the minimal prime ideals of $R[It,t^{-1}]$ are in bijection with those of $R$ (see the general discussion in [HS06] before Thm. 5.1.4), we may assume that $R$ is a domain. As $\operatorname{Frac}(R[It,t^{-1}])=\operatorname{Frac}(R)(t)$ has transcendence degree 1 over $\operatorname{Frac}(R)$, [Jaf60, Ch. I, Thm. 2, p. 10] implies that if $v^{\prime}$ is any valuation of $\operatorname{Frac}(R[It,t^{-1}])$ extending a valuation $v$ of $\operatorname{Frac}(R)$ then the rank $\operatorname{rank}(v^{\prime})$ of $v^{\prime}$ is at most $\operatorname{rank}(v)+1$. It follows that $\operatorname{vdim}R[It,t^{-1}]\leq\operatorname{vdim}R+1$. The kernel of the surjection $R[It,t^{-1}]\to R,t\mapsto 1$, is a non-trivial prime ideal, hence $\operatorname{vdim}R[It,t^{-1}]>\operatorname{vdim}R$, for example by Lemma 6.1. ∎
Proposition 6.4.
Let $R$ be a local ring and $I\subseteq R$ be a finitely generated, proper ideal with analytic spread $\ell(I)$. Then there exists an $n\geq 1$ and a reduction $J$ of $I^{n}$ which is generated by $\ell(I)$ elements.
Proof.
We follow the proof of [HS06, Prop. 8.3.8]. Let $\mathfrak{m}$ denote the maximal ideal of $R$, $\kappa=R/\mathfrak{m}$. By graded Noether normalization (see e.g. [HS06, Thm. 4.2.3]) there exists an integer $n$ and homogeneous elements $\bar{a}_{1},\dots,\bar{a}_{\ell}\in\mathcal{F}_{I}(R)_{n}=I^{n}/\mathfrak{m}I^{n}$ such that $A=\kappa[\bar{a}_{1},\dots,\bar{a}_{\ell}]\subseteq\mathcal{F}_{I}(R)$ is a polynomial ring and the extension $A\subseteq\mathcal{F}_{I}(R)$ is finite. In particular, $\ell=\dim\mathcal{F}_{I}(R)=\ell(I)$. Moreover, also the subextension $A\subseteq\bigoplus_{k\geq 0}I^{nk}/\mathfrak{m}I^{nk}$ is then finite (up to grading, the latter algebra is $\mathcal{F}_{I^{n}}(R)$). Let $nd$ be the maximal degree of a finite set of homogeneous generators. Then $I^{n(d+1)}/\mathfrak{m}I^{n(d+1)}=\sum_{i=1}^{\ell}\bar{a}_{i}\cdot I^{nd}/\mathfrak{m}I^{nd}$. Choose lifts $a_{i}\in I^{n}$ of the $\bar{a}_{i}$ and let $J=(a_{1},\dots,a_{\ell})\subseteq I^{n}$. The previous equality implies $JI^{nd}+\mathfrak{m}I^{n(d+1)}=I^{n(d+1)}$. As $I$ is finitely generated, the Nakayama lemma gives $JI^{nd}=I^{n(d+1)}$, i.e. $J$ is a reduction of $I^{n}$. ∎
6.3. Weibel vanishing